OpenStudy (anonymous):

The vertex of a parabola is (3,2). A second point on the parabola is (1,5). Which of the following points is also on the graph? (3,7) (5,5) (3,-2) (-1,5)

OpenStudy (jdoe0001):

if we use the vertex form of a parabola, that is http://nayak2010.files.wordpress.com/2011/01/picture-of-vertex-form-of-parabola.png we would end up with \(\bf y = a(x-3)^2+2\) now, what's is "a"? well, we dunno but we know that another point in the parabola is (1, 5), so let's use that in our "vertex form" equation \(\bf y = a(x-3)^2+2\qquad \qquad (1,5)\\ \quad \\ 1=a(5-3)^2+2\implies -1=4a\implies -\cfrac{1}{4}=a\\ \quad \\ \quad \\ \textit{so our parabola equation is}\qquad y = -\cfrac{1}{4}(x-3)^2+2\) now check the points given, see which "x" produces the accompanying "y" in the pair :)

OpenStudy (anonymous):

If b^p =N, then which of the following statements is also true? \[\log_{b}N = p \] \[\log_{p}N = b\] \[\log_{n} b = p\] \[N^\frac{ 1 }{ b } = p\] Please hurry and help.